# LLM-based NLG Evaluation: Current Status and Challenges

### v2

@@ 4.4 Fine-tuning Implementation
The fine-tuning process is uniformly implemented by different works on their selected open-source LLMs, like LLaMA, and respective constructed data, with some targeted settings. Specifically, Prometheus maintains balanced data distributions during fine-tuning, including the length and label. JudgeLM eliminates potential biases by randomly swapping sample pairs to be compared and randomly removing references. INSTRUCTSCORE utilizes GPT-4 to provide error annotations for the intermediate outputs of the fine-tuned model for further supervised reinforcement. And based on some preliminary experiments and manual analysis, TIGERScore determines appropriate ratios of different types of data during fine-tuning, which are claimed to be crucial by them. Moreover, CritiqueLLM implements separately, with and without references, and explores the effects of data and model scale. Compared to the vanilla fine-tuning setting, these methods have improved the efficiency of model training and the robustness of evaluations.
@@ 4.5 Pros and Cons
Those shortcomings of prompting LLM for NLG evaluation can be significantly alleviated due to the customized implementation of data construction and model fine-tuning here. For instance, most fine-tuned models range between 7B and 13B in the scale of parameters, facilitating low-cost inference use and good reproducibility, with performance close to GPT4 in NLG evaluation. And specific measures can be adopted to prevent related biases found in GPT4 during different stages. Furthermore, this type of approach allows for continuous iteration and improvement of the model to address potential deficiencies or emerging issues discovered in future applications.
However, some biases associated with GPT4 may still persist, as the data construction of most methods employs GPT4 for critical evaluation annotation. On the other hand, the base open-source LLMs selected by existing works are primarily the series of LLaMA. With the rapid updates and improvements of open-source large models recently, it adheres to the intuition that employing a more powerful base LLM should lead to better evaluation performance. However, this means repetitive fine-tuning processes and computational expenses from scratch since directly migrating existing fine-tuned models to the new base LLM is difficult.
Additionally, although many existing methods aspire for more flexible and comprehensive evaluation through fine-tuning, demanding excessive evaluation settings may ultimately lead to poor performance or failure in model training, as found by AUTO-J and CritiqueLLM on criteria and references, respectively. However, there are some disagreements here since Prometheus and JudgeLM show different results. Moreover, considering the different evaluation settings in existing works, it is challenging to conduct a horizontal comparison among them. These issues require further exploration in future research.
@@ 5 Human-LLM Collaborative Evaluation
5.1 Overview
While LLMs demonstrate robust evaluation capabilities, there exists a need for further enhancement in terms of their reliability, particularly in establishing a stronger correlation with human evaluation outcomes. Although human evaluation is the gold-standard evaluation approach in NLG, it is recognized for its associated high costs and susceptibility to subjective biases van der Lee et al. (2021); Deriu et al. (2021); Li et al. (2023b). The robust and comprehensive capabilities exhibited by LLMs underscore considerable potential for the development of collaborative evaluation methodologies that integrate both human and LLMs. In recent investigations, researchers have initiated the exploration of collaborative evaluation paradigms, which include traditional NLG evaluation methods such as scoring and explaining Zhang et al. (2021); Li et al. (2023b), broader evaluation methods such as testing and debugging Ribeiro and Lundberg (2022), and auditing NLG models to ensure fairness Rastogi et al. (2023). Furthermore, scholars Saunders et al. (2022) are actively engaging in efforts to address the intricate challenge of scalable oversight Amodei et al. (2016) through the collaboration of humans and LLMs. The objective is to devise strategies for effectively evaluating models in tasks that pose inherent difficulties for human assessors. This collaborative approach seeks to leverage the distinctive strengths of both human evaluators and sophisticated language models to achieve robust and nuanced evaluations in challenging domains.
5.2 Scoring and Explaining
Automated evaluation frequently exhibits a limited correlation with human judgments, while human evaluation, though reliable, is labor-intensive. Zhang et al. (2021) present a human-machine collaborative framework (HMCEval) which conceptualizes dialogue evaluation as a sample assignment problem to ensure the reliability of evaluation outcomes while minimizing human effort and achieves 99% accuracy with half human effort. Recently, LLMs have emerged as a cost-effective alternative to human evaluations. However, both humans and LLMs have limitations, including inherent subjectivity and unreliable judgments, especially in open-ended tasks with diverse requirements.
To address challenges associated with inconsistent evaluation criteria in open-ended tasks and explore synergy between humans and LLM-based evaluators, Li et al. (2023b) proposes a Collaborative Evaluation pipeline (COEVAL), which involves designing a checklist of task-specific criteria and conducting detailed evaluations where LLMs generate initial ideation and humans engage in scrutiny. Depending solely on score predictions is insufficient for ensuring reliable evaluation and error detection, particularly when specific criteria demand nuanced analysis beyond straightforward scoring. Building upon recent developments in explainable NLP (Yin and Neubig (2022); Jung et al. (2022); Ribeiro and Lundberg (2022); Ye et al. (2023b)), COEVAL is assigned the additional task of generating explanations to elucidate evaluation outcomes to facilitate a trustworthy collaborative evaluation process. Results indicate COEVAL effectively evaluates lengthy texts by utilizing LLMs, saving significant time and reducing human evaluation outliers. Despite the involvement of LLMs, human scrutiny remains essential, contributing to the revision of around 20% of LLM evaluation scores for enhanced reliability.
5.3 Broader Evaluation Tasks
The broader evaluation of NLG models involves testing and debugging the models. Current methods often rely on highly variable human creativity and extensive manual effort or are limited to addressing a very specific class of bugs. Ribeiro and Lundberg (2022) introduce AdaTest, a process that uses LLMs in collaboration with human feedback to automatically generate unit tests that highlight bugs in a target model, which proves to make users 5-10 times more effective at identifying bugs and assists users in effectively fixing bugs without introducing new ones. Moreover, LLMs have shown biases and irresponsible behavior, necessitating thorough auditing before deployment Blodgett et al. (2020); Jones and Steinhardt (2022). AdaTest++ Rastogi et al. (2023) draw on insights from literature on human-AI collaboration and sensemaking, and engage with research experts in safe and fair AI, which emphasizes the importance of sensemaking and effective communication between humans and AI to capitalize on their complementary strengths in collaborative auditing. AdaTest++ successfully leverages human strengths, such as schematization and hypothesis testing. Moreover, users identified a range of failure modes across 26 different topics in issues that were revealed in formal audits and those that were previously under-reported. Additionally, ensuring trustworthiness in LLMs for challenging tasks Chen et al. (2021); Nakano et al. (2021); Li et al. (2022); Menick et al. (2022) poses a crucial challenge. Scalable oversight Amodei et al. (2016) aims to effectively evaluate models on tasks challenging for humans and suggests the use of AI for assistance. Saunders et al. (2022) explored providing critiques of model outputs as a form of assistance, demonstrating that model-generated critiques assist humans in identifying overlooked flaws.
@@ 5.4 Pros and Cons
The advantages of human-AI collaborative evaluation lie in achieving a balance between efficiency and cost, as demonstrated by COEVAL Li et al. (2023b) achieving this equilibrium. Additionally, there are complementary strengths between humans and AI. For instance, AdaTest++ Rastogi et al. (2023) empowers users to consistently utilize their strengths throughout the auditing process, benefiting significantly from LLM. Users who generate the most topics heavily rely on LLM suggestions while employing their contextual reasoning and semantic understanding to update their mental models vigilantly and identify model failures.
However, there are drawbacks. The evaluation results of LLMs may be sensitive to the formats used to query the model and might require additional support for prompt writing Li et al. (2023b); Rastogi et al. (2023). Furthermore, the current capability to assess confidence levels is not strong enough, making it challenging to determine when to trust the LLM. Furthermore, certain level of human supervision is still necessary, making it less convenient and cost-effective compared to fully automated evaluation.
@@ 6 Conclusions and Future Trends
Through the above review of studies on NLG evaluation based on LLMs, we find that these four categories of approaches have their respective strengths and weaknesses, and most of the existing work is concentrated on prompting LLMs. In view of this, we offer some suggestions for future directions in this field.
@@
NLG evaluation for low-resource languages and new task scenarios. Almost all existing research focuses on English data. However, it is doubtful whether LLMs have similar levels of NLG evaluation capability for texts in other languages, especially low-resource languages. As Zhang et al. (2023a) points out, we should be more cautious about using LLMs to evaluate texts in non-Latin languages. Additionally, existing research mainly focuses on more traditional NLG tasks such as translation, summarization, and dialogue. However, there are many new scenarios in reality with different requirements and evaluation criteria. Research on low-resource languages and new task scenarios will provide a more comprehensive understanding of LLMs’ evaluation capabilities.

### v3

@@ 4.3 Fine-tuning Implementation
The fine-tuning process is implemented by different studies on their selected open-source foundation LLMs, like LLaMA, and respective constructed data, with some targeted settings. Specifically, Prometheus maintains balanced data distributions during fine-tuning, including the length and label. JudgeLM eliminates potential biases by randomly swapping sample pairs to be compared and randomly removing references. INSTRUCTSCORE utilizes GPT-4 to provide error annotations for the intermediate outputs of the fine-tuned model for further supervised reinforcement. And based on some preliminary experiments and manual analysis, TIGERScore determines appropriate ratios of different types of data during fine-tuning, which are claimed to be crucial by them. Moreover, CritiqueLLM implements separately, with and without references, and explores the effects of data and model scale. Themis employs additional rating-guided preference optimization after the fine-tuning process. Specifically, Self-Taught utilizes the evaluation results of the fine-tuned model itself for self-iterative optimization, leading to surprising improvements. Compared to the vanilla fine-tuning setting, these methods have improved the efficiency of model training and the robustness of evaluations.
@@ 4.4 Pros and Cons
The shortcomings of prompting LLMs for NLG evaluation can be significantly alleviated by the customized construction of training data and specifically fine-tuned LLMs. For instance, most models in Table 2 have less than 14B parameters, facilitating low-cost inference in practice and good reproducibility, with performance comparable to GPT4. And specific measures can be adopted to prevent certain biases found in GPT4 during different stages, such as randomly changing the order of training pairs for position bias. Furthermore, this type of approach allows for continuous iteration and improvement of the model to address potential deficiencies or emerging issues discovered in future applications.
However, some inherent biases associated with GPT4 may still persist, like self-biases, as the data construction of most methods employs GPT4 for critical evaluation annotation. On the other hand, many studies have chosen open-source foundation LLMs spanning three generations of the Llama series. With the recent rapid updates and improvements of open-source LLMs, it is intuitive that employing a more powerful foundation LLM should lead to better evaluation performance of the fine-tuned model. However, this means repetitive fine-tuning processes and computational expenses from scratch since directly migrating existing fine-tuned models to the new foundation LLM is challenging.
Additionally, although many existing methods aspire to more flexible and comprehensive evaluation through fine-tuning, demanding excessive evaluation settings may ultimately lead to poor performance or failure in model training, as AUTO-J and CritiqueLLM were found to have difficulties with criteria and references, respectively. However, there are some disagreements here since Prometheus, JudgeLM, and CompassJudger show different results, indicating such a seemingly straightforward fine-tuning process is actually quite complex. Moreover, considering the different evaluation settings in existing works, conducting a horizontal comparison among them is challenging. These issues require further exploration in future research.
@@ 5 Human-LLM Collaborative Evaluation
Human evaluation remains the gold standard for NLG due to its ability to capture nuanced aspects of quality. However, it is expensive, time-consuming, and prone to subjective biases (van der Lee et al., 2021; Deriu et al., 2021; Li et al., 2023a). While LLMs offer a promising avenue for automated evaluation, their reliability and correlation with human judgment are still areas of active development (Li et al., 2023b; Liu et al., 2023b). Human-LLM collaborative evaluation seeks to leverage the strengths of both: the nuanced judgment of humans and the scalability and efficiency of LLMs. This chapter explores emerging paradigms in this collaborative space, focusing on how humans and LLMs can work together to improve the accuracy, efficiency, and trustworthiness of NLG evaluation. This includes collaborative approaches like: traditional evaluation tasks such as scoring and explaining (Zhang, Ren, and de Rijke, 2021; Li et al., 2023a); general evaluation tasks such as testing and debugging (Ribeiro and Lundberg, 2022); auditing NLG models to ensure fairness (Rastogi et al., 2023); aligning LLM-assisted evaluation of LLM outputs with human preferences (Shankar et al., 2024); addressing the intricate challenge of scalable oversight (Raji and Dobbe, 2024; Saunders et al., 2022).
5.1 Human-Guided LLM Evaluation
Some works (Zhang, Ren, and de Rijke, 2021; Li et al., 2023a) focus on approaches where LLMs perform the primary evaluation task, but with significant guidance and oversight from humans. This guidance can take several forms, from designing detailed evaluation criteria to refining LLM outputs.
One common method is called checklist-based evaluation. A key challenge in open-ended NLG tasks is the lack of consistent evaluation criteria. Li et al. (2023a) address this with COEVAL, a collaborative pipeline where humans design a task-specific checklist. LLMs then use this checklist to generate initial evaluations and explanations, drawing on developments in explainable NLP (Yin and Neubig, 2022; Jung et al., 2022; Ribeiro and Lundberg, 2022; Ye et al., 2023). Humans then scrutinize these LLM-generated evaluations, refining scores and explanations. This approach leverages the LLM’s ability to process large amounts of text while retaining human oversight to ensure accuracy and reduce outliers. Notably, human review still leads to revisions in approximately 20% of LLM scores, highlighting the importance of human judgment. Furthermore, InteractEval (Chu, Kim, and Yi, 2025) combines human and LLM-generated attributes using Think Aloud methods to create questions and produce final prediction scores. Think Aloud methods mean that human experts verbalize their thoughts and LLMs articulate their knowledge to generate text attribute insights using sample texts and evaluation rubrics, which highlights the necessity of effectively combining humans and LLMs in an automated checklist-based text evaluation.
Collaborative assignment is also useful for human-guided LLM evaluation. Zhang, Ren, and de Rijke (2021) propose HMCEval, a framework that frames dialogue evaluation as a sample assignment problem. This approach aims to optimize the allocation of evaluation tasks between humans and machines to maximize accuracy while minimizing human effort. HMCEval achieves high accuracy (99%) with significantly reduced human involvement (half the effort). Besides, EvalAssist (Ashktorab et al., 2024) can help practitioners refine evaluation criteria using both direct and pairwise assessment strategies. Ashktorab et al. (2024) also examine how users refine their criteria and identify key differences between the two evaluation approaches examined how users refine their criteria and identified key differences between the two evaluation approaches.
5.2 LLM-Assisted Human Evaluation
Some works (Ribeiro and Lundberg, 2022; Rastogi et al., 2023; Pozdniakov et al., 2024) explore scenarios where humans remain the primary evaluators, but LLMs provide assistance to improve efficiency, identify flaws, or audit for biases.
Ribeiro and Lundberg (2022) introduce AdaTest, a system where LLMs generate unit tests to identify bugs in a target NLG model. Human feedback guides the LLM, significantly increasing the effectiveness of bug detection (5-10x improvement). This demonstrates the power of LLMs in generating diverse test cases, guided by human intuition. In the task of evaluating machine translation systems, Zouhar, Kocmi, and Sachan (2024) assist annotators by pre-filling error annotations with recall-oriented automatic quality estimation, which achieves the effect of reducing the time per span annotation by half while maintaining the same annotation quality level and further cutting the annotation budget by almost 25%.
Addressing biases and irresponsible behavior in LLMs is crucial (Blodgett et al., 2020; Jones and Steinhardt, 2022). AdaTest++ (Rastogi et al., 2023), drawing on human-AI collaboration research, facilitates collaborative auditing. Humans leverage their strengths in schematization and hypothesis testing, while LLMs assist in identifying a wide range of failure modes. This collaborative approach uncovered both previously known and under-reported issues.
Evaluating LLMs on complex tasks can be challenging even for humans (Chen et al., 2021; Nakano et al., 2021; Li et al., 2022; Menick et al., 2022). The concept of scalable oversight (Raji and Dobbe, 2024) suggests using AI to assist in evaluation. Saunders et al. (2022) explore using LLM-generated critiques to help humans identify flaws in model outputs, demonstrating that this form of assistance improves human performance. What’s more, Pozdniakov et al. (2024) focus on designing conversational user interfaces, which helps educators to use LLMs to evaluate assignments of students.
@@ 5.3 Pros and Cons
Human-LLM collaborative evaluation offers a compelling balance between the accuracy of human judgment and the efficiency of automated methods. Key advantages include: (1) Efficiency and Cost-Effectiveness: LLMs can significantly reduce the time and resources required for evaluation. (2) Complementary Strengths: Humans excel at nuanced judgment and critical thinking, while LLMs excel at processing large amounts of data and generating diverse outputs. (3) Improved Accuracy: Combining human and LLM strengths can lead to more accurate and reliable evaluations than either approach alone.
However, challenges remain: (1) Prompt Sensitivity: LLM evaluation results can be sensitive to the phrasing of prompts, requiring careful prompt engineering (Li et al., 2023a; Rastogi et al., 2023). (2) Confidence Calibration: LLMs’ ability to accurately assess their own confidence is still limited, making it difficult to know when to trust their judgments. (3) Need for Human Oversight: While reduced, human supervision is still necessary, limiting the potential for full automation. (4) Explainability: Ensuring the collaborative process is transparent and understandable can be challenging.
Method Parameter Coherence Consistency Fluency Relevance Overall
Traditional Metrics
BERTScore 355M 0.285 0.151 0.186 0.302 0.231
BARTScore 400M 0.474 0.266 0.258 0.318 0.329
LLM-derived Metrics
GPTScore (FT5) 11B 0.456 0.438 0.424 0.343 0.415
GPTScore (OPT) 66B 0.359 0.453 0.380 0.337 0.382
GPTScore (GPT-3) 175B 0.434 0.449 0.403 0.381 0.417
GPTScore (Phi-4) 14B 0.319 0.436 0.386 0.154 0.324
GPTScore (Llama-3.1) 70B 0.415 0.478 0.437 0.288 0.405
GPTScore (Qwen-2.5) 72B 0.447 0.486 0.437 0.376 0.436
Prompting LLMs
G-Eval (GPT-3.5) - 0.440 0.386 0.424 0.385 0.409
G-Eval (GPT-4) - 0.582 0.507 0.455 0.548 0.523
Phi-4 14B 0.479 0.454 0.421 0.452 0.451
Llama-3.1 70B 0.510 0.387 0.317 0.494 0.427
Qwen-2.5 72B 0.515 0.509 0.435 0.528 0.497
Fine-tuning LLMs
INSTRUCTSCORE 7B 0.328 0.232 0.260 0.211 0.258
Prometheus 2 7B 0.403 0.318 0.269 0.356 0.336
Themis 8B 0.566 0.600 0.571 0.474 0.553
TIGERScore 13B 0.381 0.427 0.363 0.366 0.384
CompassJudger-1 32B 0.494 0.424 0.318 0.410 0.411
Human-LLM Collaborative Evaluation
InteractEval (GPT-3.5 1st) - 0.583 0.630 0.734 0.614 0.640
InteractEval (GPT-3.5 2nd) - 0.590 0.614 0.726 0.623 0.638
InteractEval (GPT-4 1st) - 0.649 0.799 0.783 0.626 0.714
InteractEval (GPT-4 2nd) - 0.660 0.781 0.816 0.642 0.725
Table 4: Performance of different types of LLM-based NLG evaluation approaches on SummEval, where some results are from Fu et al. (2024), Hu et al. (2024) and Chu, Kim, and Yi (2025).
@@ 6 Conclusions and Future Trends
6.1 Comparison with traditional evaluation metrics.
Traditional evaluation metrics are criticized for their poor correlation with human judgments (Stent, Marge, and Singhai, 2005), uninterpretable evaluation results (Zhang, Vogel, and Waibel, 2004), and inability to adapt to specific evaluation criteria (Wiseman, Shieber, and Rush, 2017), which are being greatly mitigated by LLM-based evaluation. However, the higher cost, the requirements for computing resources, and the issues of reproducibility may be the downside.
6.2 Comparison between different types of LLM-based NLG evaluation.
We compare different types of LLM-based evaluation according to flexibility and reproducibility due to the difficulty of comparing the effectiveness of different types of methods in various scenarios.
Flexibility
Human-LLM Collaborative Evaluation > Prompting LLMs > Fine-tuning LLMs > LLM-derived Metrics. Human-LLM Collaborative Evaluation involves human annotators, which provides the highest flexibility. LLM-derived Metrics are typically designed to evaluate specific aspects, such as text similarity, and do not fully allow evaluation criteria to be expressed in natural language, making them the least flexible. When comparing Prompting LLMs and Fine-tuning LLMs, the former, which uses proprietary models, generally performs better at following instructions compared to smaller open-source models.
Reproducibility
LLM-derived Metrics $\approx$ Prompting LLMs > Fine-tuning LLMs > Human-LLM Collaborative Evaluation. Human-LLM Collaborative Evaluation requires human annotators, and the recruitment and training of these annotators pose greater challenges to reproducibility. LLM-derived Metrics and Prompting LLMs do not modify the existing models, and therefore have better reproducibility than Fine-tuning LLMs. However, they may still become non-reproducible if proprietary models are deprecated.
Performance
Human-LLM Collaborative Evaluation > Fine-tuning LLMs $\approx$ Prompting LLMs > LLM-derived Metrics. We compare the performance of different LLM-based evaluation approaches on the most commonly used NLG evaluation benchmark on summarization, SummEval (Fabbri et al., 2021), as shown in Table 4. When using the same LLMs, LLM-derived metrics perform worse than directly prompting LLMs for evaluation, and the latter is more convenient. Moreover, among methods of fine-tuning LLMs, only models focused on NLG evaluation scenarios, such as Themis, outperform prompting-based methods, including that using GPT-4. Other studies either use relatively outdated foundation LLMs or lack training on specific evaluation aspects like those in SummEval, leading to relatively weaker performance. Furthermore, Human-LLM Collaborative Evaluation enhances the LLM evaluation by incorporating checklists elaborated with human expert insights and LLM knowledge, resulting in the strongest performance.
Cost
LLM-derived Metrics $\approx$ Prompting open-source LLMs < Fine-tuning LLMs $\approx$ Prompting proprietary LLMs < Human-LLM Collaborative Evaluation. When using the same open-source LLM, the inference costs of LLM-derived metrics, prompting LLM, and fine-tuning LLM methods are the same, while fine-tuning LLM incurs additional training costs. When prompting proprietary LLMs, the cost is high and mainly concentrated in API calls during evaluation, making it difficult to directly compare with the training cost required for fine-tuning LLM. Moreover, human-LLM collaborative evaluation requires the involvement of human experts for each task, making it the most expensive approach.
6.3 Future Directions
@@
NLG evaluation for low-resource languages and new task scenarios. Almost all existing research focuses on English data. However, it is doubtful whether LLMs have similar levels of NLG evaluation capability for texts in other languages, especially low-resource languages. As (Zhang et al., 2024) points out, we should be more cautious about using LLMs to evaluate texts in non-Latin languages. We believe that the lack of evaluation capability of LLM-based evaluators on low-resource languages may be due to the insufficient presence of these languages in the pretraining corpus. Therefore, further fine-tuning on certain low-resource languages may be a potential strategy to address this issue, and Hada et al. (2024a) have already shown promising preliminary results. Additionally, existing research mainly focuses on more traditional NLG tasks such as translation, summarization, and dialogue. However, there are many new scenarios in reality with different requirements and evaluation criteria. For example, using LLMs to automatically evaluate scientific reviews could be valuable in identifying and flagging content that is unfaithful or unclear, alerting reviewers to potential issues. Research on low-resource languages and new task scenarios will provide a more comprehensive understanding of LLMs’ evaluation capabilities.
@@
Acknowledgements.
This work was supported by Beijing Science and Technology Program (Z231100007423011) and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology). We appreciate the anonymous reviewers for their helpful comments. Xiaojun Wan is the corresponding author.
[…]